# Real-time Bayesian inference at extreme scale:

### v1

@@ IV Current State of the Art
Unfortunately, our operator is not low rank, due to the hyperbolic nature of the forward wave propagation problem (which tends to preserve information), and the location of the pressure sensors directly on the seafloor, whose motion we seek to infer. In our numerical experiments on smaller but representative regions [henneking2025goal], the effective rank is nearly of the order of the data dimension, and thus we can expect CG to take $\mathcal{O}$(250,000) iterations. Since each iteration requires the action of $\mathbf{F}$ and $\mathbf{F}^{*}$ on a vector, i.e., a pair of forward/adjoint wave propagations, we would expect CG to take $\mathcal{O}$(50) years (on 512 GPUs) for our target problem. And this is for computing $\mathbf{m}_{\text{map}}$ via (2); in the absence of low-rank properties, the even more daunting task of evaluating the uncertainty of the inverse solution (which is also needed for the posterior of the QoI, $\mathbf{\Gamma}_{\hskip-2.0pt\text{post}(q)}$) via the inverse of the Hessian in (3) is unthinkable. While approximations of high (global) rank Hessians have been attempted, they have not been viable for high-frequency wave problems in $10^{9}$ parameter dimensions, e.g., [ambartsumyan2020hierarchical].
@@
Alternatively, we could relax the need to exactly solve (2) for $\mathbf{m}_{\text{map}}$ and instead seek a surrogate approximation of the p2o map, $\mathbf{F}$. However, this too is a monumental challenge: constructing surrogates in $10^{9}$ parameter dimensions, where each training point amounts to a forward wave propagation ($\sim$1 hour on 512 GPUs), is essentially impossible using conventional surrogate methods. Specially-architected neural operator surrogates that exploit the geometry of p2o maps have been developed, scaled to high ambient dimensions, and deployed within Bayesian inverse problems, e.g., [CaoOLearyRoseberryGhattas25]. However, the efficiency of these surrogates (measured by number of training data needed) is predicated on low-rank representations of $\mathbf{F}$, again a property not enjoyed by our p2o maps. Thus, surrogates are not a viable option for our problem.
@@
Finally, we might attempt to construct a projection-based reduced order model (ROM) of the forward acoustic–gravity wave equations, by projecting (1) onto a low-dimensional reduced basis. The ROM would have to be parameterized over $10^{9}$ inputs, which is well beyond the reach of current methods. A more fundamental difficulty is that efficient ROMs for high-frequency wave propagation are not viable due to the Kolmogorov $N$-width problem [greif2019kolmogorov], which asserts the non-existence of a low-dimensional subspace embedding.
@@
Next, we present novel parallel algorithms that exploit the autonomous nature of the dynamics (1) and p2o and p2q maps $\mathbf{F}$ and $\mathbf{F}_{\hskip-1.0ptq}$ to induce an offline–online decomposition permitting real-time solution of both the Bayesian inverse problem for seafloor motion and subsequent wave height forecasting.
@@ V Innovations Realized
Our real-time Bayesian inference and prediction framework combines the following key ideas: (i) an offline–online decomposition of the inverse solution that precomputes various mappings to enable the real-time application of the inverse operator and entirely circumvents the need for PDE solutions during the inversion phase; (ii) representation of the inverse operator in data space instead of high-dimensional parameter space; (iii) extension of this real-time inference methodology to the goal-oriented setting for the prediction of QoI under uncertainty; (iv) exploitation of the shift invariance of (1) to reduce the required number of PDE solutions by several orders of magnitude; and (v) exploitation of this same property to perform efficient FFT-based and multi-GPU-accelerated Hessian matrix–vector products (matvecs). We employ this framework to construct a digital twin for tsunami early warning in the CSZ that solves a Bayesian inverse problem with over $10^{9}$ parameters and 8,820 QoI exactly (up to rounding errors); both the parameter inference and the complete end-to-end, data-to-inference-to-prediction computations can be carried out within a fraction of a second.
Figure 3: Physics-based magnitude 8.7 dynamic rupture earthquake scenario for a margin-wide rupture in the CSZ, from left to right: (a) True seafloor displacement; (b) snapshot of true seafloor acoustic pressure field with 600 hypothesized sensor locations; (c) snapshot of true sea surface wave height; (d) inferred mean of seafloor displacement; (e) uncertainties plotted as pointwise standard deviations in meters of seafloor normal displacement; and (f) snapshot of reconstructed sea surface wave height with 21 locations for QoI predictions. Hyperlinks to animations of: (i) source (seafloor vertical uplift and normal velocity); (ii) forward solution (sea bottom pressure and surface wave height); and (iii) inverse solution (true and inferred seafloor normal displacement).
@@ V-A FFT-based Hessian matvecs
Solving the inverse problem (2) requires numerous Hessian matvecs, and each Hessian matvec comes at the cost of a pair of forward and adjoint PDE solutions. For large-scale problems, PDE solutions are computationally expensive and cannot be performed in real time. One key innovation that enables real-time inference is to recognize and exploit the fact that the p2o map is governed by an autonomous dynamical system. Specifically, the acoustic–gravity model (1) represents a linear time-invariant (LTI) dynamical system for which the PDE operator does not explicitly depend on time.
The LTI dynamical system structure of the governing equations implies that the discrete p2o map $\mathbf{F}$, which maps parameters $\mathbf{m}$ (spatiotemporal seafloor motion) to observables $\mathbf{d}$ (sea bottom pressure) via solution of the discretized equations of the acoustic–gravity PDE model (1), inherits a special structure. Let $N_{m}$ denote the spatial parameter points, $N_{d}$ the number of sensors ($N_{d}\ll N_{m}$), and $N_{t}$ the temporal dimension of parameters and observations [^1] For simplicity of presentation, without loss of generality, we assume that the frequency of the observations and QoI predictions is equal to the temporal discretization of the parameters, i.e., $N_{t}=N_{t}^{d}=N_{t}^{q}$. ($N_{t}\gg 1$), i.e.,
@@
In other words, the block Toeplitz structure of $\mathbf{F}$ corresponds to a time-shift invariance of its blocks $\mathbf{F}_{ij}$. This shift invariance implies that (i) the p2o map can be precomputed from only $N_{d}$ (number of sensors) adjoint PDE solutions (corresponding to the first block column of $\mathbf{F}$) and stored compactly at cost of $\mathcal{O}(N_{m}N_{d}N_{t})$ memory; (ii) the block Toeplitz matrix can be embedded within a block circulant matrix that is block-diagonalized by the discrete Fourier transform [gray2006toeplitz]; and (iii) the p2o matvec then becomes a block-diagonal matvec operation in Fourier space.
These FFT-based p2o matvecs can be implemented efficiently on multi-GPU clusters. In particular, care is taken to arrange the data layouts of the matrices and vectors in each part of the computation to minimize the amount of GPU–GPU communication. By exchanging the order of space and time vector indices local to each processor, the need for strided memory accesses during the core computational kernels is avoided. A 2D GPU partitioning scheme is employed, where the dimensions of the processor grid are adaptively tuned according to the problem sizes and total number of GPUs in order to further reduce communication costs [venkat2024fft].
@@
Finally, the entire discussion about the p2o map $\mathbf{F}$ analogously applies to the p2q map $\mathbf{F}_{\hskip-1.0ptq}\in\mathbb{R}^{(N_{q}N_{t})\times(N_{m}N_{t})}$ that maps the parameters $\mathbf{m}$ to the QoI $\mathbf{q}\in\mathbb{R}^{N_{q}N_{t}}$ (sea surface wave heights) at $N_{q}$ forecast locations ($N_{q}\ll N_{m}$).
@@ V-B Real-time Bayesian inference and prediction framework
To enable real-time Bayesian inference, our framework decomposes the inverse solution into several precomputation (offline) Phases 1–3 and a real-time inference (online) Phase 4 [henneking2025goal], as depicted in Fig. 2.
@@
Phase 1 performs $N_{d}+N_{q}$ (number of sensors plus QoI forecasts) adjoint PDE solutions to precompute the block Toeplitz matrices $\mathbf{F}$ and $\mathbf{F}_{\hskip-1.0ptq}$.
@@
Phase 2 computes the block Toeplitz matrices $\{\mathbf{G}^{*},\mathbf{G}_{q}^{*}\}=\mathbf{\Gamma}_{\hskip-2.0pt\text{prior}}\{\mathbf{F}^{*},\mathbf{F}_{\hskip-1.0ptq}^{*}\}$ by performing $N_{d}+N_{q}$ solves of the inverse elliptic operator defining the prior covariance $\mathbf{\Gamma}_{\hskip-2.0pt\text{prior}}$. Then, the Sherman–Morrison–Woodbury formula is used to rewrite the posterior covariance as
@@
$\mathbf{\Gamma}_{\hskip-2.0pt\text{post}}=\left(\mathbf{F}^{*}\mathbf{\Gamma}_{\hskip-2.0pt\text{noise}}^{-1}\mathbf{F}+\mathbf{\Gamma}_{\hskip-2.0pt\text{prior}}^{-1}\right)^{-1}\!=\left(\mathbf{\Gamma}_{\hskip-2.0pt\text{prior}}-\mathbf{G}^{*}\mathbf{K}^{-1}\mathbf{G}\right),$
@@
where $\mathbf{K}=\mathbf{\Gamma}_{\hskip-2.0pt\text{noise}}+\mathbf{F}\mathbf{G}^{*}$. Doing so effectively shifts the inverse operator from the intractable high-dimensional parameter space (of dimension $N_{m}N_{t}$) to the still-large-but-tractable data space (of dimension $N_{d}N_{t}$); for that reason, we refer to $\mathbf{K}$ as the (prior-preconditioned) “data space Hessian.” The dense $\mathbf{K}$ matrix is precomputed with $N_{d}N_{t}$ matvecs on unit vectors, each of which—thanks to the FFT-based matvecs of $\mathbf{F}$ and $\mathbf{G}^{*}$—can be executed within fractions of a second, after which $\mathbf{K}$ can be Cholesky-factorized.
@@
Phase 3 computes the uncertainties for the QoI predictions by forming the dense covariance matrix $\mathbf{\Gamma}_{\hskip-2.0pt\text{post}(q)}=\mathbf{F}_{\hskip-1.0ptq}\mathbf{\Gamma}_{\hskip-2.0pt\text{post}}\mathbf{F}_{\hskip-1.0ptq}^{*}$. This requires $N_{q}N_{t}$ matvecs with $\mathbf{\Gamma}_{\hskip-2.0pt\text{post}(q)}$ on unit vectors; each matvec effectively solves a billion-parameter inverse problem (in the actions of $\mathbf{\Gamma}_{\hskip-2.0pt\text{post}}$ on a vector), which is made tractable thanks to our ability to solve the inverse problem in a fraction of a second. Analogously, we precompute the data-to-QoI map, $\mathbf{Q}\coloneq\mathbf{F}_{\hskip-1.0ptq}\mathbf{\Gamma}_{\hskip-2.0pt\text{post}}\mathbf{F}^{*}\mathbf{\Gamma}_{\hskip-2.0pt\text{noise}}^{-1}$, which, for a moderate number of QoI forecasts, can (in Phase 4) be used to make rapid real-time QoI predictions directly from observational data, thereby bypassing the need to explicitly reconstruct the seafloor velocity parameters (see §VII).
@@
Phase 4 executes the (online) parameter inference and QoI prediction, which, given data $\mathbf{d}^{\text{obs}}$, reduces to computing the MAP points, $\mathbf{m}_{\text{map}}=\mathbf{\Gamma}_{\hskip-2.0pt\text{post}}\mathbf{F}^{*}\mathbf{\Gamma}_{\hskip-2.0pt\text{noise}}^{-1}\mathbf{d}^{\text{obs}}$ and $\mathbf{q}_{\text{map}}=\mathbf{F}_{\hskip-1.0ptq}\mathbf{m}_{\text{map}}=\mathbf{Q}\mathbf{d}^{\text{obs}}$ (taking $\mathbf{m}_{\text{prior}}=\boldsymbol{0}$), both of which, after the precomputations done in Phases 1–3, can be executed in real time.
@@ V-C Application to the Cascadia subduction zone
To assess the feasibility of our framework for near-field real-time tsunami forecasting in Cascadia without assuming simplified earthquake models, we generated seafloor displacements that capture the nonlinear interaction of seismic wave propagation and frictional rock failure during a realistic CSZ rupture scenario computed with a 3D dynamic earthquake rupture model [glehman2024partial, uphoff2017sumatra]. We then used these “true” seafloor displacements as a source for our acoustic–gravity forward model (i.e. normal displacements $b(\vec{x},t)$ in (1)) to generate synthetic observational data with 1% relative added noise, at $N_{d}=600$ hypothesized seafloor pressure sensors. Using these sparse, noisy observations of acoustic pressure, we then solved the PDE-based inverse problem for the spatiotemporal parameter field representing seafloor velocities during the rupture event, and made predictions of the QoI (sea surface wave height) at $N_{q}=21$ forecast locations. We note that even a small number of sea surface wave height forecasts is informative for early warning purposes [liu2021comparison].
[…]

### v2

@@ IV Current State of the Art
Unfortunately, our operator is not low rank, due to the hyperbolic nature of the forward wave propagation problem (which tends to preserve information), and the location of the pressure sensors directly on the seafloor, whose motion we seek to infer. In our numerical experiments on smaller but representative regions [21], the effective rank is nearly of the order of the data dimension, and thus we can expect CG to take $\mathcal{O}$(250,000) iterations. Since each iteration requires the action of $\mathbf{F}$ and $\mathbf{F}^{*}$ on a vector, i.e., a pair of forward/adjoint wave propagations, we would expect CG to take $\mathcal{O}$(50) years (on 512 GPUs) for our target problem. And this is for computing $\mathbf{m}_{\text{map}}$ via (2); in the absence of low-rank properties, the even more daunting task of evaluating the uncertainty of the inverse solution (which is also needed for the posterior of the QoI, $\mathbf{\Gamma}_{\thinspace\text{post}(q)}$) via the inverse of the Hessian in (3) is unthinkable. While approximations of high (global) rank Hessians have been attempted, they have not been viable for high-frequency wave problems in $10^{9}$ parameter dimensions, e.g., [22].
@@
Alternatively, we could relax the need to exactly solve (2) for $\mathbf{m}_{\text{map}}$ and instead seek a surrogate approximation of the p2o map, $\mathbf{F}$. However, this too is a monumental challenge: constructing surrogates in $10^{9}$ parameter dimensions, where each training point amounts to a forward wave propagation ($\sim$1 hour on 512 GPUs), is essentially impossible using conventional surrogate methods. Specially-architected neural operator surrogates that exploit the geometry of p2o maps have been developed, scaled to high ambient dimensions, and deployed within Bayesian inverse problems, e.g., [23]. However, the efficiency of these surrogates (measured by number of training data needed) is predicated on low-rank representations of $\mathbf{F}$, again a property not enjoyed by our p2o maps. Thus, surrogates are not a viable option for our problem.
@@
Finally, we might attempt to construct a projection-based reduced order model (ROM) of the forward acoustic–gravity wave equations, by projecting (1) onto a low-dimensional reduced basis. The ROM would have to be parameterized over $10^{9}$ inputs, which is well beyond the reach of current methods. A more fundamental difficulty is that efficient ROMs for high-frequency wave propagation are not viable due to the Kolmogorov $N$-width problem [24], which asserts the non-existence of an accurate low-dimensional subspace embedding.
@@
Next, we present novel parallel algorithms that exploit the autonomous nature of the dynamics (1) and p2o and p2q maps $\mathbf{F}$ and $\mathbf{F}_{\thinspace q}$ to induce an offline–online decomposition permitting real-time solution of both the Bayesian inverse problem for seafloor motion and subsequent wave height forecasting.
@@ V Innovations Realized
Our real-time Bayesian inference and prediction framework combines the following key ideas: (i) an offline–online decomposition of the inverse solution that precomputes various mappings to enable the real-time application of the inverse operator and entirely circumvents the need for PDE solutions during the inversion phase; (ii) representation of the inverse operator in data space instead of high-dimensional parameter space; (iii) extension of this real-time inference methodology to the goal-oriented setting for the prediction of QoI under uncertainty; (iv) exploitation of the shift invariance of (1) to reduce the required number of offline PDE solutions by several orders of magnitude; and (v) exploitation of this same property to perform efficient FFT-based and multi-GPU-accelerated Hessian matrix–vector products (matvecs). We employ this framework to construct a digital twin for tsunami early warning in the CSZ that solves a Bayesian inverse problem with over $10^{9}$ parameters and 8,820 QoI exactly (up to rounding errors); both the parameter inference and the complete end-to-end, data-to-inference-to-prediction computations can be carried out within a fraction of a second.
Figure 3: Physics-based magnitude 8.7 dynamic rupture earthquake scenario for a margin-wide rupture in the CSZ, from left to right: (a) true seafloor displacement; (b) snapshot of true seafloor acoustic pressure field with 600 hypothesized sensor locations; (c) snapshot of true sea surface wave height; (d) inferred mean of seafloor displacement; (e) uncertainties plotted as pointwise standard deviations in meters of seafloor normal displacement; and (f) snapshot of reconstructed sea surface wave height with 21 locations for QoI predictions. Hyperlinks to animations of: (i) source (seafloor vertical uplift and normal velocity); (ii) forward solution (seafloor pressure and surface wave height); and (iii) inverse solution (true and inferred seafloor normal displacement).
@@ V-A FFT-based Hessian matvecs
Solving the inverse problem (2) requires numerous Hessian matvecs, each of which comes at the cost of a pair of forward/adjoint PDE solutions. For large-scale problems, PDE solutions are computationally expensive and cannot be performed in real time. One key innovation that enables real-time inference is to recognize and exploit the fact that the p2o map is governed by an autonomous dynamical system. Specifically, the acoustic–gravity model (1) represents a linear time-invariant (LTI) dynamical system for which the PDE operator does not explicitly depend on time.
The LTI dynamical system structure of the governing equations implies that the discrete p2o map $\mathbf{F}$, which maps parameters $\mathbf{m}$ (spatiotemporal seafloor velocity) to observables $\mathbf{d}$ (seafloor pressure) via solution of the discretized equations of the acoustic–gravity PDE model (1), inherits a special structure. Let $N_{m}$ denote the spatial parameter points, $N_{d}$ the number of sensors ($N_{d}\ll N_{m}$), and $N_{t}$ the temporal dimension of parameters and observations [^1] For simplicity of presentation, and without loss of generality, we assume that the temporal discretization of the parameters is taken the same as the frequency of the observations and QoI predictions, i.e., $N_{t}=N_{t}^{d}=N_{t}^{q}$. ($N_{t}\gg 1$), i.e.,
@@
In other words, the block Toeplitz structure of $\mathbf{F}$ corresponds to a time-shift invariance of its blocks $\mathbf{F}_{ij}$. This shift invariance implies that (i) the p2o map can be precomputed from only $N_{d}$ (number of sensors) adjoint PDE solutions (corresponding to the first block column of $\mathbf{F}$) and stored compactly at cost of $\mathcal{O}(N_{m}N_{d}N_{t})$ memory; (ii) the block Toeplitz matrix can be embedded within a block circulant matrix that is block-diagonalized by the discrete Fourier transform [25]; and (iii) the p2o matvec then becomes a block-diagonal matvec operation in Fourier space.
These FFT-based p2o matvecs can be implemented efficiently on multi-GPU clusters. In particular, care is taken to arrange the data layouts of the matrices and vectors in each part of the computation to minimize the amount of GPU–GPU communication. By exchanging the order of space and time vector indices local to each processor, the need for strided memory accesses during the core computational kernels is avoided. A 2D GPU partitioning scheme is employed, where the dimensions of the processor grid are adaptively tuned according to the problem sizes and total number of GPUs in order to further reduce communication costs [26].
@@
Finally, the entire discussion about the p2o map $\mathbf{F}$ analogously applies to the p2q map $\mathbf{F}_{\thinspace q}\in\mathbb{R}^{(N_{q}N_{t})\times(N_{m}N_{t})}$ that maps the parameters $\mathbf{m}$ to the QoI $\mathbf{q}\in\mathbb{R}^{N_{q}N_{t}}$ (sea surface wave heights) at $N_{q}$ forecast locations ($N_{q}\ll N_{m}$).
@@ V-B Real-time Bayesian inference and prediction framework
To enable real-time Bayesian inference, our framework decomposes the inverse solution into several precomputation (offline) Phases 1–3 and a real-time inference (online) Phase 4 [21], as depicted in Fig. 2.
@@
Phase 1 performs $N_{d}+N_{q}$ (number of sensors plus QoI forecasts) adjoint PDE solutions to precompute the block Toeplitz matrices $\mathbf{F}$ and $\mathbf{F}_{\thinspace q}$.
@@
Phase 2 computes the block Toeplitz matrices $\{\mathbf{G}^{*},\mathbf{G}_{q}^{*}\}=\mathbf{\Gamma}_{\thinspace\text{prior}}\{\mathbf{F}^{*},\mathbf{F}_{\thinspace q}^{*}\}$ by performing $N_{d}+N_{q}$ solves of the inverse elliptic operator defining the prior covariance $\mathbf{\Gamma}_{\thinspace\text{prior}}$. Then, the Sherman–Morrison–Woodbury formula is used to rewrite the posterior covariance as
@@
$\mathbf{\Gamma}_{\thinspace\text{post}}=\left(\mathbf{F}^{*}\mathbf{\Gamma}_{\thinspace\text{noise}}^{-1}\mathbf{F}+\mathbf{\Gamma}_{\thinspace\text{prior}}^{-1}\right)^{-1}\!=\left(\mathbf{\Gamma}_{\thinspace\text{prior}}-\mathbf{G}^{*}\mathbf{K}^{-1}\mathbf{G}\right),$
@@
where $\mathbf{K}=\mathbf{\Gamma}_{\thinspace\text{noise}}+\mathbf{F}\mathbf{G}^{*}$. Doing so effectively shifts the inverse operator from the intractable high-dimensional parameter space (of dimension $N_{m}N_{t}$) to the still-large-but-tractable data space (of dimension $N_{d}N_{t}$); for that reason, we refer to $\mathbf{K}$ as the (prior-preconditioned) “data space Hessian.” The dense $\mathbf{K}$ matrix is precomputed with $N_{d}N_{t}$ matvecs on unit vectors, each of which—thanks to the FFT-based matvecs of $\mathbf{F}$ and $\mathbf{G}^{*}$—can be executed within fractions of a second, after which $\mathbf{K}$ can be Cholesky-factorized.
@@
Phase 3 computes the uncertainties for the QoI predictions by forming the dense covariance matrix $\mathbf{\Gamma}_{\thinspace\text{post}(q)}=\mathbf{F}_{\thinspace q}\mathbf{\Gamma}_{\thinspace\text{post}}\mathbf{F}_{\thinspace q}^{*}$. This requires $N_{q}N_{t}$ matvecs with $\mathbf{\Gamma}_{\thinspace\text{post}(q)}$ on unit vectors; each matvec effectively solves a billion-parameter inverse problem (in the actions of $\mathbf{\Gamma}_{\thinspace\text{post}}$ on a vector), which is made tractable thanks to our ability to solve the inverse problem in a fraction of a second. Analogously, we precompute the data-to-QoI map, $\mathbf{Q}\coloneq\mathbf{F}_{\thinspace q}\mathbf{\Gamma}_{\thinspace\text{post}}\mathbf{F}^{*}\mathbf{\Gamma}_{\thinspace\text{noise}}^{-1}$, which, for a moderate number of QoI forecasts, can (in Phase 4) be used to make rapid real-time QoI predictions directly from observational data, thereby bypassing the need to explicitly reconstruct the seafloor velocity parameters (see §VII).
@@
Phase 4 executes the (online) parameter inference and QoI prediction, which, given data $\mathbf{d}^{\text{obs}}$, reduces to computing the MAP points, $\mathbf{m}_{\text{map}}=\mathbf{\Gamma}_{\thinspace\text{post}}\mathbf{F}^{*}\mathbf{\Gamma}_{\thinspace\text{noise}}^{-1}\mathbf{d}^{\text{obs}}$ and $\mathbf{q}_{\text{map}}=\mathbf{F}_{\thinspace q}\mathbf{m}_{\text{map}}=\mathbf{Q}\mathbf{d}^{\text{obs}}$ (taking $\mathbf{m}_{\text{prior}}=\boldsymbol{0}$), both of which, after the precomputations done in Phases 1–3, can be executed in real time.
@@ V-C Application to the Cascadia subduction zone
To assess the feasibility of our framework for near-field real-time tsunami forecasting in Cascadia without assuming simplified earthquake models, we generated seafloor displacements that capture the nonlinear interaction of seismic wave propagation and frictional rock failure during a realistic CSZ rupture scenario computed with a 3D dynamic earthquake rupture model [27, 4]. We then used these “true” seafloor displacements as a source for our acoustic–gravity forward model (i.e. normal displacements $b(\vec{x},t)$ in (1)) to generate synthetic observational data with 1% relative added noise, at $N_{d}=600$ hypothesized seafloor pressure sensors. Using these sparse, noisy observations of acoustic pressure, we then solved the PDE-based inverse problem for the spatiotemporal parameter field representing seafloor velocities during the rupture event, and made predictions of the QoI (sea surface wave height) at $N_{q}=21$ forecast locations. We note that even a small number of sea surface wave height forecasts is informative for early warning purposes [28].
To solve this inverse problem, we discretized the acoustic–gravity model (1) with high-order finite elements in MFEM (see §VI-B) using a hexahedral mesh with a spatial resolution of 300 meters (less in shallow areas of the CSZ) with 3.74 billion states and explicit Runge–Kutta time-stepping, with a timestep size dictated by the CFL condition, for a simulation time of 420 seconds. The parameter field was discretized with $N_{m}=$ 2,416,530 spatial points and $N_{t}=$ 420 temporal steps (1 Hz frequency) for a total parameter dimension of $\sim$1.015 billion. Results of the parameter inference and QoI prediction under uncertainty are shown in Figs. 3 and 4; note that the inverse solution is depicted in terms of the seafloor normal displacement, $\int_{0}^{T}m(\vec{x},t)dt$, which is visually simpler to interpret than snapshots of the inferred spatiotemporal parameter field (seafloor velocities). The computation times and other performance-related aspects for this inverse problem are presented in §VII.